# In Search of Cosmic Time: Complete Observables and the Clock Hypothesis

### v1

@@ 4. Hubble Flow and the Clock Hypothesis
For a given matter model we can then write the density and pressure in terms of the scale factor allowing us to for example re-write the first equation as:
$H^{2}(\tau)=H^{2}_{0}\left[\Omega_{m,0}\frac{1}{a(\tau)^{3}}+\Omega_{r,0}\frac{1}{a(\tau)^{4}}+\Omega_{\Lambda}+\Omega_{k}\frac{1}{a(\tau)^{2}}\right]$
@@
The question is then whether Hubble parameter $H(\tau)$ and the proper time along the Hubble flow can be understood to be measurable quantities. Let us consider each in turn.
$d_{L}(z)=(1+z)\int_{0}^{z}\frac{c\,dz^{\prime}}{H(z^{\prime})}$
By measuring the redshift $z$ through spectrometers and the distance $d_{L}$ through large-field telescopes, such as the Hubble Space Telescope, we can obtain $H(z)$, from which the value of the Hubble parameter at different times $\tau$ can be derived. In particular, since $\tau$ is related to the redshift through the scale factor $a(\tau)=1/(1+z)$, we can thus obtain information about $H(\tau)$.
Second, and more subtle, is the question of whether we can measure the proper time along the Hubble flow $\tau$. Again we evidently cannot measure directly the proper time of our galaxy following the Hubble flow. If we say that we use whatever periodic physical system as a clock ‘attached to the galaxy’, it will not follow the Hubble flow. Actually, the concept of Hubble flow can be valid only at cosmological scales, so even for our Galaxy we should account for the effects of peculiar velocities. Completely eliminating peculiar motions from measurements of galaxy recession velocities is not possible, but it is possible to correct them in an approximate way. Therefore, strictly speaking, experimentally we cannot measure directly with a clock the proper time of any object following the Hubble flow. What we can do is appeal to the clock hypothesis. This amounts to the hypothetical assumption there is a clock that measures proper time along any given world-line. Proper time can be rigorously defined in the context of a relativistic spacetime $(M,g_{ab})$ as follows (Malament 2012, §2.3): let $\gamma:[s_{1},s_{2}]\to M$ be a smooth future-directed timelike curve in the manifold $M$ with tangent $\xi^{a}$. Then the proper time associated with the curve relative to the metric $g_{ab}$ is given by:
[…]
s where $ds$ is the line element. Since the clock hypothesis applies also to the world-lines of observers following the Hubble flow, we seem to have solved the problem by stipulation.
@@
Since we can associate measurements with partial observables and partial observables with reference frames, we can wonder: what kind of Reference Frame the cosmic proper time is? And it is a bona-fide partial observable? Given the clock hypothesis, we can stipulate a clock that measures cosmic time. Furthermore, cosmic time will always be dynamically coupled with the expansion rate of the Universe, parametrised by $H(\tau)$, since both quantities depend on the same FLRW metric. The cosmic time value, being a proper time, is a structural property of the gravitational field given by equation (7). This is very similar to the the sense in which the Hubble parameter measures the rate of expansion of the volume of the universe as a geometric quantity, which is derived from the gravitational field: $dV=\sqrt{-g}dx^{\mu}$. $\tau$ is a geodesic reference clock, since the four-velocity of the cosmological fluid is associated to the geodesic dynamics of a dust fluid, whose energy-momentum tensor $T_{ab}=\rho U_{a}U_{b}$ is source of the EFEs and give rise to the FLRW solution. Thus, $\tau$ is a RRF in (Bamonti 2023)’s classification. The nature of the RRF clock comes from the fact that $\tau$ is the proper time of the cosmological fluid, whose back-reaction on gravity is taken into account and gives rise to the FLRW metric which in turn determines the proper time $\tau$ (this is the essence of the non-linear feedback of EFEs). This RRF clock provides the privileged representation in which the cosmic microwave background radiation is represented as perfectly homogeneous and isotropic, in absence of small inhomogeneities of the primordial universe.
We thus have that cosmic proper time: i) is an RRF since it involved back-reaction; ii) is a bona-fide partial observable according to Bamonti and Gomes (2024); and iii) corresponds to a measurable quantity by the clock hypothesis. Is this enough for us to conclude that $H(\tau)$ is a complete observable and thus have solved Thiemann’s dilemma? Almost.
@@
As stated above, the problem is that on a practical level, it is not possible to have an experimentally accessible clock, (i.e. with which we can exchange signals), that follows the Hubble flow. In general, distant galaxies are considered to follow the Hubble Flow, as for very distant galaxies, the contribution of their peculiar motions is negligible compared to the recession velocity due to the expansion of the universe. Thus, let’s consider a galaxy in our past light cone as a satellite sending radiation towards us. In this way, we would use it as a kind of Rovelli (2002a)’s ‘GPS clock that would allow us to define local quantities, such as the Hubble parameter, in its proper time. [^15] Note also that clusters of matter represent inhomogeneities that are assumed to evolve following the underlying FLRW background structure. So, their evolution does not influence the global FLRW evolution. “More precisely, it is assumed that effects from the small scale inhomogeneities onto the largest scales can be neglected, i.e. there is no substantial backreaction” Schander and Thiemann (2021). Thus, galaxies, clusters and other agglomerates of matter are treated as test particles. Again, however, the cosmic time would be the time measured by some stipulated and not further defined clock ‘attached’ to that galaxy-satellite and broadcasting the measured value to the experimenter, via light signals. Apart from the experimental problem to construct a valuable experimental setting, and the need to take into account the galaxy’s peculiar velocities, there remains the problem of defining an origin of such time measured by the clock-galaxy. One solution would be to conveniently place the origin as the ‘zero time’ of the formation event of such a cluster. Or also, analogous to the construction of GPS coordinates in Rovelli (2002a), as the time of the galaxy’s encounter with another galaxy. In any case, the proper time of the galaxy will never be a global time. As Rovelli states: ‘Our Galaxy and Andromeda are heading towards a collision: when they will meet, the times elapsed from the Big Bang will be different in the two galaxies. None of the two will have any claim of being more of a “true” time than the other.’ (Rovelli 2024, p.18). It is practically impossible to have a clock measuring the proper time parametrising the Hubble flow.
In this context it is interesting to note that Brown and Read (2021) argue that the clock hypothesis is not strictly realised and is better understood as ‘clock condition’. In a similar vein, one might view the clock hypothesis as constituting the definition of clocks as objects that measure proper time. The important point is that if the clock hypothesis holds, then we are able to treat the $\tau$ in the Friedmann equations as a measurable quantity just like the Hubble parameter. However, in the context of cosmology and the Hubble flow, there is a tension between applying the clock hypothesis and the idea of a clock as a real, experimentally accessible physical system. Real physical clocks do not seem to be the kind of things that can in fact measure proper time along the Hubble flow. In the following section we return to the ideas of partial and complete observable and RRFs to better understand both this challenge and the comparative merits of the approach of Thiemann described in the previous section.
@@ 5. Finding Cosmic Time
Let us recap. One the one hand, the widely used and accepted criterion for an observable in a theory with temporal diffeomorphism symmetry is that such observables should be Dirac observables and therefore have (weakly) vanishing Poisson bracket with all first class constraints. On the other hand, the widely used and empirically established Friedmann equations describing the dynamics of the scale factor can be understood to correspond to those generated by a first class constraint in a theory with temporal diffeomorphism symmetry. It seems like we must either give up on the Dirac criterion for observables or modify our understanding cosmological dynamics.
@@
We have considered two alternative responses to this dilemma as follows. First, Thiemann argues that we should adopt the second option and demonstrates how we might explicitly reconstruct Friedmann cosmology as a de-parameterized theory based upon a phantom matter field acting as a physical clock that measures cosmic time. A deviation parameter then quantifies how much the dynamics of the universe, governed by the modified Friedmann equation, differs from the standard model of cosmology. Second, we have constructed an alternative approach that re-interprets Friedmann equations as evolution equations parameterized by proper time, rather than coordinate time. On this approach we understand the equations as describing the correlation between two independently measurable ‘partial observables’ given by the Hubble parameter and proper time along the Hubble flow. The Hubble parameter is a measurable quantity within modern cosmology. Furthermore, following the clock hypothesis, we have that since clocks measure proper time along world-lines, a clock following the Hubble flow will necessarily measure cosmic time. The problem, however, is that the proper time along the Hubble flow does not correspond to a physical quantity associated with a measuring procedure by a clock leading to a number, and so it seems we are no longer implementing the partial and complete observables procedure in the spirit of Rovelli (2002b). The crucial point is to define an experimental measure of cosmic time consistent with the clock hypothesis.
In this context, it is worth noting again that Thiemann holds a different understanding of the partial and complete observables approach to Rovelli. In particular, for Thiemann partial ‘observables’ are not observables at all and so there is not sense in which they need to be associate to a measuring procedure. We can thus understand the Thiemann approach to complete observables and cosmic time as built upon abandoning not one by two conventionally accepted aspects of the formalism, viz. the clock hypothesis and the distinction between partial and complete observability. An observable simpliciter is defined as a Dirac Observable and there is no requirement that such observables are built out of independently measurable functions, which however are not (Dirac) observables. On this way of thinking, $H(\tau)$ will not be an observable simpliciter, and should not be expected to commute with the Hamiltonian constraint.
@@
Moreover, since the clock hypothesis is abandoned, $\tau$ is not observable, and for that $H(\tau)$ is not a (Dirac) observable too. Plausibly, it is precisely the abandonment of the clock hypothesis which led Thiemann to use a phantom scalar clock as the physical, observable clock of the theory. In any case, the crucial point is that adopting this perspective does not amount simply to an alternative interpretation of the theory. Rather, it is to reformulate the classical theory of cosmology such in a way that modifies the observable consequences. What modifications are made will depend upon the clock choice. However, there is no choice that corresponds to a strict preservation of the Friedmann equations: ‘whatever matter is used for deparametrisation, there will be corrections […] This should have observable consequences!’ (Thiemann 2006, p.9). Non-standard observable consequences are of course a virtue in a physical theory. Furthermore it is worth noting that Thiemann’s approach in the paper in question is connected to a specific research programme in terms of Phantom k-essence cosmology (Aguirregabiria et al. 2004) and thus the modifications in question could be independently motivated and in principle tested via the relevant modified matter or gravity theory. There are thus good methodological reasons to pursue more radical approach to identifying cosmic time.
@@
Can we plot a plausible path towards a more conservative approach that preserves both the distinct notions of partial and complete observable and the clock hypothesis? Recall once more that on the original Rovelli definition a partial observable is a physical quantity associated with measuring procedure leading to a number. This definition fits well with the way the Hubble parameter features in our cosmological observational practice, even if its measurement is indirect. The problem was that proper time along the Hubble flow is not associated with a direct measuring procedure in any straightforward sense that involves a clock. Notwithstanding the clock hypothesis, there seems to be a tension between our intuitive notion of measurement of time (by a clock) and the role played by cosmic proper time in our scientific theories. It is worth considering at this juncture, however, that the intuitive notion of measurement has itself been disputed in the context of a practice orientated account of scientific measurement. This has lead to a reorientation of the philosophy of scientific measurement towards a model-based account. Furthermore, perhaps the most detailed and powerful example of the conceptual heavy lifting that a model-based account of measurement can do is in context of the measurement of time. We will see that the path, similarly to the case of $H$, will be to rely on indirect measurements and not measurements made by a clock, as is commonly done when dealing with time measurements.
@@
Let us briefly consider key ideas from the pathbreaking work of Tal (2016) on the measurement of time via atomic clocks cf. Thébault (2021a). Following this account we recognise that the time ‘measured’ by atomic clocks does not correspond to a simple procedure of reading a number from a device. Rather, ‘Coordinated Universal Time’ (UTC) is based upon a standardisation procedure involving multiple atomic clocks distributed throughout the globe and systematic modelling at various stages. Caesium plays a particularly important role in modern time-keeping since it is transitions of an idealised caesium atom that are the basis for the definition of the second. However, as emphasised by Tal, this does not mean that one can simply read seconds from real caesium atoms. The caesium atom that defines the second is an idealised construct, at rest at zero degrees Kelvin and with no coupling to any external fields. Actual atomic clocks are then built to approximately realise the ideal caesium clock, with known sources of difference minimised and modelled. However, the ‘primary’ standards caesium clocks typically only operates for a few weeks at a time in order to calibrate ‘secondary standards’. The secondary standards are a different class of atomic clocks that are less accurate but can be run continuously for a number of years. The secondary standard clocks also must be modelled. In particular, the ‘readings’ of the clocks are subject to quantitive adjustments relating to the known sources of difference between their ideal physical operation and their actual physical realisation. This allows the time that they read to be a close approximation to that read by their idealised counterpart. The crucial point is that UTC is not ‘read’ by either primary or secondary standards. Rather it is a product of a further abstraction based upon the readings of the different participant clocks throughout the world. Furthermore, not only are different clocks weighted differently in UTC, since some clocks are more noisy, but since the clocks are at different physical locations on the earth, one must also take account of their differing proper times, as determined by the relevant differences in gravitational field and four acceleration (Tal 2016, p. 302). The general implication for of a model-based account of measurement for the definition of a partial observable is thus to substantially liberalise the sense of ‘procedure’ in Rovelli’s definition. In particular, on a model-based account of measurement, the procedure involved may include not just indirect measurement but various tears of modelling, calibration, and aggregation. Furthermore, and more importantly for our discussion, comparison between the status of UTC and cosmic time throws into relief our failure in the latter case to find an actual physical system that plays the role of the clock. There are no clocks that measure UTC either.
@@
Consider, then, that we can ‘measure’ cosmic time via various cosmological phenomena. For example, we can use the analysis of the power spectrum of temperature fluctuations of the CMB (TT-power spectrum) measured by the PLANCK satellite. In particular, the CMB power spectrum represents the power distribution of the temperature anisotropies as a function of angular scale. The different angular scales correspond to the scales of the baryonic acoustic oscillations (BAOs) which are pressure waves generated by interactions between radiation (photons) and matter (baryons) in the primordial plasma before decoupling era. Anisotropies reflect primordial density differences, which led to the formation of galaxies and other structures (Kolb 2018, ch.9). The power spectrum contains peaks and valleys at different scales and their position and amplitude are sensitive to cosmological parameters. For example, a higher density of matter leads to higher peaks that are closer together, and a flat universe tends to have a different distribution of peaks than a curved universe. The Hubble constant $H_{0}$ also affects the scale of the fluctuations and the position of the peaks. [^16] Formally, the TT-power spectrum is the expansion in Legendre polynomials $\mathfrak{P}$ of the variance of temperature fluctuations: $\langle\cfrac{\Delta T}{T}(\gamma_{1})\cfrac{\Delta T}{T}(\gamma_{2})\rangle=1/{2\pi}\sum_{l}{l(l+1)C_{l}\mathfrak{P}_{l}(cos\theta)},$ where $\gamma$ represents the direction in which temperature fluctuations $\Delta T$ are measured and $C_{l}$ is the coefficient of the $l$-th multipole, representing the amplitude of temperature fluctuations at a certain angular scale $\theta\propto 1/l$. There is a privileged angular scale for these oscillations, which corresponds to the size of the sound horizon at the time of recombination. This angular scale corresponds approximately to the multipole $l=200$ with a dependence on cosmological parameters. See Figure 1.
Figure 1. Temperature power spectra for three different Hubble constants. The red curve represents the model with $H_{0}=70$ Km/s/MPc. The blue curve represents the model with $H_{0}=100$ Km/s/MPc. The green curve represents the model with $H_{0}=30$ Km/s/MPc. For the acquisition of the data to be plotted, we used the CAMB web interface https://lambda.gsfc.nasa.gov/toolbox/camb_online.html.
@@
By reconstructing the power spectrum it is possible to determine cosmological time. In particular, cosmological time is determined by integrating the equations of the expansion of the universe:
$\tau(z)=\int_{z}^{\infty}\frac{dz^{\prime}}{(1+z^{\prime})H(z^{\prime})},$

### v2

@@ 4. Hubble Flow and the Clock Hypothesis
where we have used the fact that $\frac{\ddot{a}}{a}=\dot{H}+H^{2}$ and assumed differentiation with respect to proper time. For a given matter model we can then write the density and pressure in terms of the scale factor allowing us to for example re-write the first equation as:
(13) $H^{2}(\tau)=H^{2}_{0}\left[\Omega_{m,0}\frac{1}{a(\tau)^{3}}+\Omega_{r,0}\frac{1}{a(\tau)^{4}}+\Omega_{\Lambda}+\Omega_{k}\frac{1}{a(\tau)^{2}}\right],$
where $H_{0}$ is the Hubble constant and we have introduced experimentally measurable density parameters at the current time; $\Omega_{0,R}$ for the radiation density, $\Omega_{0,M}$ for the matter (dark plus baryonic) density, $\Omega_{0,k}$ for the spatial curvature density, and $\Omega_{0,\Lambda}$ for the cosmological constant density. [^21] The measurement of $H_{0}$ suffers from the so-called Hubble-tension. For a review see Smeenk (2022). The question is then whether and in what sense Hubble parameter $H(\tau)$ and the proper time along the Hubble flow can be understood to be measurable quantities. Let us consider each in turn.
(14) $d_{L}(z)=(1+z)\int_{0}^{z}\frac{c\,dz^{\prime}}{H(z^{\prime})}$
By measuring the redshift $z$ through spectrometers and the distance $d_{L}$ through large-field telescopes, such as the Hubble Space Telescope, we can infer $H(z)$. However, notice that this procedure requires ‘fitting’ noisy discrete data, introducing mild model dependence. From $H(z)$ the value of the Hubble parameter $H(\tau)$ at different times $\tau$ can be derived. In fact, using the definition of redshift in terms of the scale factor $a$, namely $a(\tau)=\frac{1}{1+z}$, and the fact that $H=\frac{\dot{a}}{a}$, we can derive the following expression:
(15) $H(z)=-\frac{\dot{z}}{1+z}.$
This equation highlights that, since it is impractical to directly measure $\dot{z}$, we must infer its behaviour using a specific cosmological model. The construction of $H(\tau)$ requires combining observations with a theoretical model that relates the cosmic time $\tau$ to the redshift $z$ [^22] See Section 5, formula (17), where a way to calculate $\tau(z)$ is provided. Thus, any inference of $H$ from redshift measurements likely involves some degree of model dependence. The procedure to transition from $H(z)$ to $H(\tau)$ can be summarised as follows: I) Infer $H(z)$ using formula (14); II) Compute $\tau(z)$ according to the chosen theoretical model (see Section 5); III) Invert the relation $\tau(z)$ to obtain $z(\tau)$; IV) Combine $H(z)$ with $z(\tau)$ to construct $H(\tau)$. This procedure underscores the fact that $H$ and $\tau$ are not completely independent: knowledge of one requires knowledge of the other. Their (dynamical) interdependence supports the claim that $H(\tau)$ is a complete observable, in line with the remark outlined in Section 2, proposed by Bamonti and Gomes (2024). For completeness, we note that model dependence is not unique to the measurement of $H(z)$. Many cosmological parameters rely on similar assumptions.
Second, and more subtle, is the question of whether we can measure the proper time along the Hubble flow. Again we evidently cannot measure directly the proper time $\tau$ of our galaxy following the Hubble flow. If we say that we use whatever periodic physical system as a clock ‘attached to the galaxy’, it will not follow the Hubble flow. Actually, the concept of Hubble flow can be valid only at cosmological scales, so even for our Galaxy we should account for the effects of peculiar velocities. Completely eliminating peculiar motions from measurements of galaxy recession velocities is not possible, but it is possible to correct them in an approximate way. Therefore, strictly speaking, experimentally we cannot measure directly with a clock the proper time of any object following the Hubble flow. What we can do is appeal to the clock hypothesis. This amounts to the hypothetical assumption there is a clock that measures proper time along any given world-line. Proper time can be rigorously defined in the context of a relativistic spacetime $(M,g_{ab})$ as follows (Malament 2012b, §2.3): let $\gamma:[s_{1},s_{2}]\to M$ be a smooth future-directed timelike curve in the manifold $M$ with tangent $\xi^{a}$. Then the proper time associated with the curve relative to the metric $g_{ab}$ is given by:
@@
where $ds$ is the line element. Since the clock hypothesis applies also to the world-lines of observers following the Hubble flow, we seem to have solved the problem by stipulation. Since we can associate measurements with partial observables and partial observables with reference frames, we can wonder: what kind of Reference Frame the cosmic proper time is? And it is a bona fide partial observable? Given the clock hypothesis, we can stipulate a clock that measures cosmic time. Furthermore, cosmic time will always be dynamically coupled with the expansion rate of the Universe, parametrised by $H(\tau)$, since both quantities depend on the same FLRW metric. The cosmic time value, being a proper time, is a structural property of the gravitational field given by equation (16). This is very similar to the sense in which the Hubble parameter measures the rate of expansion of the volume of the universe as a geometric quantity, which is derived from the gravitational field: $dV=\sqrt{-g}dx^{\mu}$. $\tau$ is a geodesic reference clock, since the four-velocity of the cosmological fluid is associated to the geodesic dynamics of a dust fluid, whose energy-momentum tensor $T_{ab}=\rho U_{a}U_{b}$ is source of the EFEs and give rise to the FLRW solution. Thus, $\tau$ is an RRF in (Bamonti 2023)’s classification. The nature of the RRF clock comes from the fact that $\tau$ is the proper time of the cosmological fluid, whose back-reaction on gravity is taken into account and gives rise to the FLRW metric which in turn determines the proper time $\tau$ (this is the essence of the non-linear feedback of EFEs). This RRF clock provides the privileged representation in which the cosmic microwave background radiation is represented as perfectly homogeneous and isotropic, in absence of small inhomogeneities of the primordial universe. We thus have that cosmic proper time: i) is an RRF since it involved back-reaction; ii) is a bona fide partial observable according to Bamonti and Gomes (2024); and iii) corresponds to a measurable quantity by the clock hypothesis. Is this enough for us to conclude that $H(\tau)$ is a complete observable and thus have solved Thiemann’s dilemma? Almost.
@@
As stated above, the problem is that on a practical level, it is not possible to have an experimentally accessible clock (i.e. with which we can exchange signals), that follows the Hubble flow. In general, distant galaxies are considered to follow the Hubble Flow, as for very distant galaxies, the contribution of their peculiar motions is negligible compared to the recession velocity due to the expansion of the universe. Thus, let’s consider a galaxy in our past light cone as a satellite sending radiation towards us. In this way, we would use it as a kind of Rovelli (2002a)’s ‘GPS clock’ that would allow us to define local quantities, such as the Hubble parameter, in its proper time. [^23] Note also that clusters of matter represent inhomogeneities that are assumed to evolve following the underlying FLRW background structure. So, their evolution does not influence the global FLRW evolution. “More precisely, it is assumed that effects from the small scale inhomogeneities onto the largest scales can be neglected, i.e. there is no substantial backreaction” Schander and Thiemann (2021). Thus, galaxies, clusters and other agglomerates of matter are treated as test particles. Again, however, the cosmic time would be the time measured by some stipulated and not further defined clock ‘attached’ to that galaxy-satellite and broadcasting the measured value to the experimenter, via light signals. Apart from the experimental problem to construct a valuable experimental setting, and the need to take into account the galaxy’s peculiar velocities, there remains the problem of defining an origin of such time measured by the clock-galaxy. One solution would be to conveniently place the origin as the ‘zero time’ of the formation event of such a galaxy. Or also, analogous to the construction of GPS coordinates in Rovelli (2002a), as the time of the galaxy’s encounter with another galaxy. In any case, the proper time of the galaxy will never be a global time. As Rovelli states: ‘Our Galaxy and Andromeda are heading towards a collision: when they will meet, the times elapsed from the Big Bang will be different in the two galaxies. None of the two will have any claim of being more of a “true” time than the other.’ (Rovelli 2024, p.18). It is practically impossible to have a clock measuring the proper time parametrising the Hubble flow. [^24] In a similar vein, Brown and Read (2016), note that “For any given clock, no matter how ideal its performance when inertial, there will in principle be an acceleration-producing external force, or even tidal effects inside the clock, such that the clock “breaks”, in the sense of violating the clock hypothesis. Might it not be more appropriate to call it the clock condition?” In this context, one might view the clock hypothesis as constituting the definition of clocks as objects that measure proper time. The important point is that if the clock hypothesis holds, then we are able to treat the $\tau$ in the Friedmann equations as a measurable quantity just like the Hubble parameter. However, in the context of cosmology and the Hubble flow, there is a tension between applying the clock hypothesis and the idea of a clock as a real, experimentally accessible physical system. There do not exist real physical systems that approximate clocks which measure proper time along the Hubble flow. In the following section we return to the ideas of partial and complete observable and RRFs to better understand both this challenge and the comparative merits of the approach of Thiemann described in the previous section.
@@ 5. Finding Cosmic Time
Let us recap. One the one hand, the widely used and accepted criterion for an observable in a theory with temporal diffeomorphism symmetry is that such observables should be Dirac observables and therefore have (weakly) vanishing Poisson bracket with all first class constraints. On the other hand, the widely used and empirically established Friedmann equations describing the dynamics of the scale factor can be understood to correspond to those generated by a first class constraint in a theory with temporal diffeomorphism symmetry. It seems like we must either give up on the Dirac criterion for observables or modify our understanding cosmological dynamics. We have considered two alternative responses to this dilemma as follows. First, Thiemann argues that we should adopt the second option and demonstrates how we might explicitly reconstruct Friedmann cosmology as a de-parametrised theory based upon a phantom matter field acting as a physical clock that measures cosmic time. A deviation parameter then quantifies how much the dynamics of the universe, governed by the modified Friedmann equation, differs from the standard model of cosmology. Second, we have constructed an alternative approach that re-interprets Friedmann equations as evolution equations parametrised by proper time, rather than coordinate time. On this approach we understand the equations as describing the correlation between two independently measurable ‘partial observables’ given by the Hubble parameter and proper time along the Hubble flow. The Hubble parameter is a measurable quantity within modern cosmology. Furthermore, following the clock hypothesis, we have that since clocks measure proper time along world-lines, a clock following the Hubble flow will necessarily measure cosmic time. The problem, however, is that the proper time along the Hubble flow does not correspond to a physical quantity associated with a measuring procedure by a clock leading to a number, and so it seems we are no longer implementing the partial and complete observables procedure in the spirit of Rovelli (2002b). The crucial issue is to define an experimental measure of cosmic time consistent with the clock hypothesis.
@@
In this context, it is worth noting again that Thiemann holds a different understanding of the partial and complete observables approach to Rovelli. In particular, for Thiemann partial ‘observables’ are not observables at all and so there is no sense in which they need to be associate to a measuring procedure. We can thus understand the Thiemann approach to complete observables and cosmic time as built upon abandoning not one by two conventionally accepted aspects of the formalism, viz. the clock hypothesis and the distinction between partial and complete observability. An observable simpliciter is defined as a Dirac Observable and there is no requirement that such observables are built out of independently measurable functions, which however are not (Dirac) observables. On this way of thinking, $H(\tau)$ will not be an observable simpliciter, and should not be expected to commute with the Hamiltonian constraint. In fact, since the clock hypothesis is abandoned, $\tau$ is not observable, and for that $H(\tau)$ is not a (Dirac) observable either. Plausibly, it is precisely the abandonment of the clock hypothesis which led Thiemann to use a phantom scalar clock as the physical, observable clock of the theory. In any case, the crucial point is that adopting this perspective does not amount simply to an alternative interpretation of the theory. Rather, it is to reformulate the classical theory of cosmology such in a way that modifies the empirical consequences. What modifications are made will depend upon the clock choice. However, there is no choice that corresponds to a strict preservation of the Friedmann equations: ‘whatever matter is used for deparametrisation, there will be corrections […] This should have observable consequences!’ (Thiemann 2006, p.9). Non-standard empirical consequences are of course a virtue in a physical theory. Furthermore it is worth noting that Thiemann’s approach in the paper in question is connected to a specific research programme in terms of Phantom k-essence cosmology (Aguirregabiria et al. 2004) and thus the modifications in question could be independently motivated and in principle tested via the relevant modified matter or gravity theory. There are thus good methodological reasons to pursue more radical approach to identifying cosmic time.
@@
Can we plot a plausible path towards a more conservative approach that preserves both the distinct notions of partial and complete observable and the clock hypothesis? Recall once more that on the original Rovelli’s definition a partial observable is a physical quantity associated with measuring procedure leading to a number. This definition fits well with the way the Hubble parameter features in our cosmological observational practice, even if its measurement is indirect. The problem was that proper time along the Hubble flow is not associated with a direct measuring procedure in any straightforward sense that involves a clock. Notwithstanding the clock hypothesis, there seems to be a tension between our intuitive notion of measurement of time (by a clock) and the role played by cosmic proper time in our scientific theories. It is worth considering at this juncture, however, that the intuitive notion of measurement has itself been disputed in the context of a practice orientated account of scientific measurement. This has lead to a reorientation of the philosophy of scientific measurement towards a model-based account of measurement. Furthermore, perhaps the most detailed and powerful example of the conceptual heavy lifting that a model-based account of measurement can do is in context of the measurement of time.
@@
Let us briefly consider key ideas from the path-breaking work of Tal (2016) on the measurement of time via atomic clocks cf. Thébault (2021a). Following this account we recognise that the time ‘measured’ by atomic clocks does not correspond to a simple procedure of reading a number from a device. Rather, ‘Coordinated Universal Time’ (UTC) is based upon a standardisation procedure involving multiple atomic clocks distributed throughout the globe and systematic modelling at various stages. Caesium plays a particularly important role in modern time-keeping since it is transitions of an idealised caesium atom that are the basis for the definition of the second. However, as emphasised by Tal, this does not mean that one can simply read seconds from real caesium atoms. The caesium atom that defines the second is an idealised construct, at rest at zero degrees Kelvin and with no coupling to any external fields. Actual atomic clocks are built to approximately realise the ideal caesium clock, with known sources of difference minimised and modelled. However, the ‘primary’ standards caesium clocks typically only operates for a few weeks at a time in order to calibrate ‘secondary standards’. The secondary standards are a different class of atomic clocks that are less accurate but can be run continuously for a number of years. The secondary standard clocks also must be modelled. In particular, the ‘readings’ of the clocks are subject to quantitive adjustments relating to the known sources of difference between their ideal physical operation and their actual physical realisation. This allows the time that they read to be a close approximation to that read by their idealised counterpart. The crucial point is that UTC is not ‘read’ by either primary or secondary standards. Rather it is a product of a further abstraction based upon the readings of the different participant clocks throughout the world. Furthermore, not only are different clocks weighted differently in UTC, since some clocks are more noisy, but since the clocks are at different physical locations on the earth, one must also take account of their differing proper times, as determined by the relevant differences in gravitational field and four-acceleration (Tal 2016, p. 302). One also requires a procedure for synchronisation which introduces an element of conventionality. However, despite including conventional elements, the synchronisation of the clock network towards UTC is partially anchored in underlying regularities in nature which are required for the explanation of successful stabilisation of a synchronisation standards in ‘metrological practice’ (Tal 2016, §3.1). [^25] For specific technical details regarding the relativity of synchronisation in the context of UTC and the GPS system see the discussion of Ashby (2003).
@@
The general implication for a model-based account of measurement for the definition of a partial observable is thus to substantially liberalise the sense of ‘procedure’ in Rovelli’s definition. In particular, on a model-based account of measurement, the procedure involved may include not just indirect measurement but various tiers of modelling, calibration, and aggregation. Furthermore, and more importantly for our discussion, comparison between the status of UTC and cosmic time throws into relief our failure in the latter case to find an actual physical system that plays the role of the clock. There are no actual physical systems that measure UTC either. Consider, then, that we can ‘measure’ cosmic time via various cosmological phenomena. For example, we can use the analysis of the power spectrum of temperature fluctuations of the CMB (TT-power spectrum) measured by the PLANCK satellite. In particular, the CMB power spectrum represents the power distribution of the temperature anisotropies as a function of angular scale. The different angular scales correspond to the scales of the baryonic acoustic oscillations (BAOs) which are pressure waves generated by interactions between radiation (photons) and matter (baryons) in the primordial plasma before decoupling era. Anisotropies reflect primordial density differences, which led to the formation of galaxies and other structures (Kolb and Turner 1994, ch.9). The power spectrum contains peaks and valleys at different scales and their position and amplitude are sensitive to cosmological parameters. For example, a higher density of matter leads to higher peaks that are closer together, and a flat universe tends to have a different distribution of peaks than a curved universe. The Hubble constant $H_{0}$ also affects the scale of the fluctuations and the position of the peaks.
@@
By reconstructing the power spectrum it is possible to determine cosmological time. In particular, cosmological time can be determined by integrating the equations of the expansion of the universe:
(17) $\tau(z)=\int_{z}^{\infty}\frac{dz^{\prime}}{(1+z^{\prime})H(z^{\prime})},$